### v1

Abstract
[SUMMARY] The attention mechanism has been proven to be an effective way to improve spiking neural network (SNN). However, based on the fact that the current SNN input data flow is split into tensors to process on GPUs, none of the previous works consider the properties of tensors to implement an attention module. This inspires us to rethink current SNN from the perspective of tensor-relevant theories. Using tensor decomposition, we design the projected full attention (PFA) module, which demonstrates excellent results with linearly growing parameters. Specifically, PFA is composed by the linear projection of spike tensor (LPST) module and attention map composing (AMC) module. In LPST, we start by compressing the original spike tensor into three projected tensors using a single property-preserving strategy with learnable parameters for each dimension. Then, in AMC, we exploit the inverse procedure of the tensor decomposition process to combine the three tensors into the attention map using a so-called connecting factor. To validate the effectiveness of the proposed PFA module, we integrate it into the widely used VGG and ResNet architectures for classification tasks. Our method achieves state-of-the-art performance on both static and dynamic benchmark datasets, surpassing the existing SNN models with Transformer-based and CNN-based backbones.
keywords:
[…]
1 Introduction
Spiking neural network (SNN) is currently attracting the interest of academics due to its lower energy consumption and greater bio-interpretability compared to traditional artificial neural network (ANN). The brain-inspired computation model gives it significant potential in processing temporal data. Though training a deep SNN is still a challenge, recent developments in SNN introduce backpropagation [3] to relieve this issue to some degree. Moreover, it makes the extension of ANN modules into SNN possible, such as batch normalization and residual blocks.
As it is now feasible to incorporate ANN modules into SNNs while maintaining their inherent power efficiency, we can leverage these modules to augment the overall performance of SNNs. For instance, Zheng et al. [45] propose threshold-dependent batch normalization to solve the problems of gradient vanishing and threshold-input balance. The well-known residual block proposed by He et al. [14] makes it possible to train an ultra-deep ANN network. To introduce residual block into the field of SNN, Hu et al. [17] investigate converting residual networks and propose a new technique to adjust the parameter weight correspondingly. Furthermore, Fang et al. [11] propose the SEW-ResNet to solve gradient vanishing problems and achieve identity mapping. In addition, Hu et al. [18] design several novel residual blocks and test their applicability, greatly extending the depth of a directly trained SNN. In summary, several strategies in ANN, such as normalization, architecture, and training methods, could favor the performance of SNN effectively, which also motivates us to explore more useful techniques in ANNs for better-conducting SNNs.
Figure 1: Accuracy on CIFAR10 (left) and CIFAR100 (right). Compared with other methods, PFA significantly improves the network performance.
As one of the most promising techniques in ANNs, the attention mechanism can mimic the human ability to focus on certain things while ignoring others. Recently, three representative works have proven that the attention mechanism can successfully work in SNNs. Especially, Yao et al. [42] switch channel attention to the temporal dimension, revealing that paying attention to the temporal dimension is effective for performance enhancement. Zhu et al. [48] apply two 1D convolutions to a generated 2D tensor to achieve temporal-channel attention. Inspired by experimental observations on predictive attentional remapping, Cai et al. [4] design the SCTFA module to assess the input tensor from temporal, channel, and spatial dimensions, achieving attention among these three dimensions.
Current software frameworks, i.e. PyTorch, and tensorflow, are boosted by GPU using parallel computation. This makes it hard to realize event-driven as events occur asynchronously. Based on these frameworks, as a trade-off, a popular approach is splitting the input event streams into slices and composing tensors to convenient later computation. This approach is adopted by all attention mechanism-related works, showing a promising outcome. Since we can now only efficiently process event-converted tensors, it is natural for us to consider involving tensor analysis theories to implement an attention module for SNN, which is ignored by previous work. Simultaneously, this analysis highlights the inherent resilience within SNN modules. It demonstrates that even when employing low-rank execution for Attention mechanisms, substantial performance enhancements can be achieved. This aligns perfectly with the sparsity philosophy of SNNs.
In this paper, we propose the projected-full attention (PFA), which considers projecting one input tensor into three separate lower-dimensional tensors, then calculate a full-attention map by utilizing the inverse procedure of the tensor decomposition, fully exploiting the peculiar temporal dimension in SNN with minimal computing complexity. The input tensor is first projected along three separate dimensions by the linear projection of spike tensor (LPST) module. Then, the projections are composed into an attention map by the attention map composing (AMC) module. The input tensor is fused with the attention map by the Hadamard product to obtain the refined tensor. The main contributions of this paper can be summarized as follows:
  • We propose the projected full-attention (PFA) module to achieve three-dimensional attention. Compared with previous works, PFA is a lightweight module with fewer learnable parameters, and the number of parameters grows linearly as the size of input data increases. PFA is founded on a well-established mathematical tool, thus enhancing its reliability and mathematical interpretability.
  • We present a comprehensive theoretical analysis that encompasses two key aspects: the rank of tensors and connecting factor. Based on these analyses, we propose a selection criterion for that is tailored for distinct datasets. This criterion effectively elucidates the experimental outcomes and offers valuable guidance for future applications.
  • We conduct extensive experiments on both static and dynamic datasets. Experimental results show that our PFA is very effective and achieves state-of-the-art (SOTA) accuracy on both dynamic and static datasets. In addition, ablation studies demonstrate the structural reasonableness of PFA. We visualize the attention map obtained by PFA to intuitively illustrate the attention distribution among temporal-channel-spatial dimensions.
2 Related Works and Motivation
Spiking Neural Network: Spiking neural network (SNN), the third generation of artificial neural network [29], has the potential to behave as efficiently as the human brain. In an SNN, information is transmitted through spikes, which are binary events that represent the firing of a neuron. Unlike traditional ANNs, which use continuous signals, SNN neurons communicate over time using discrete spikes. Compared with ANN, SNN is more biologically realistic by adopting models from neural dynamics and is more hardware-friendly thanks to the discrete spike. Due to its temporal property, it is supposed to have better temporal expressiveness than ANN. Quantities of studies have already been conducted. For neurodynamic models, Hodgkin and Huxley proposed the Hodgkin-Huxley (H-H) model [15] in 1952. Models like Izhikevich [19] and Leaky Integrate and Fire (LIF) [25] are suggested to strike a balance between computing complexity and bio-interpretability. Since the spiking behavior is non-differentiable, training an SNN is challenging. Several methods have been explored to address this difficulty. Some techniques [38; 39; 44] are gradient-based optimization. Others, like Rueckauer et al. [36] and Diehl et al. [9], are based on converting an ANN to SNN. There also exists ANN-SNN collaborative networks [41]. Besides the conventional datasets (static datasets) such as MNIST [7], and CIFAR10 [23], SNNs also support datasets collected from event flow cameras (dynamic datasets) like NCaltech-101 [32]. The concept behind converted datasets, of which ES-ImageNet [28], is an example to transform ordinary datasets into neural forms through specific algorithms. Another way to convert ordinary data is to recapture it using dynamic vision sensor (DVS) cameras. NCaltech-101 [32] and CIFAR10DVS [26] are in this category. As aforementioned, SNNs are hardware-friendly, making hardware implementation one of the frontier areas. Related developments include TrueNorth [2], Loihi [6], and Tianjic [33] chips. Nonetheless, there are still many obstacles in applying SNNs on various tasks, especially in its performance. In this paper, we successfully increases the classification accuracy of SNNs with the proposed PFA module.
Attention Mechanism in SNN: Since Google proposed recurrent models of visual attention in 2014 [31], the attention mechanism become a potent tool to increase the performance of a neural network, which gives distinct parts of the input data varied weights. For instance, the squeeze-and-excitation network (SENet) [16] and other attention blocks have been frequently used in traditional ANNs. From this fact, it is of vital significance to introduce the attention mechanism into SNNs. To achieve temporal-wise attention, inspired by SENet, TA-SNN [42] is proposed. Besides, TCJA [48] use two 1D convolutions to pay attention to both channel dimension and temporal dimension to prolong the attention mechanism. Nevertheless, they fail to considered three crucial aspects of SNN, i.e., temporal, channel, and spatial. By changing SENet and leaving affection in the temporal direction, Cai et al. [4] and Yao et al. [43] successfully apply temporal-channel-spatial attention to SNN. These researches reveal the fact that attention mechanism does improve the performance of SNN.
Motivation: As aforementioned, the application of the attention mechanism has demonstrated promising ability in SNN-related tasks. Regardless of the dimension(s) considered, the attention mechanism is always beneficial [42; 48; 4; 43]. Especially, the temporal-channel-spatial attention could faithfully depict the relationship of temporal, channel and spatial dimensions, thus it becomes a quite promising approach in SNNs. For this type of attention, although 3D convolution is used to accomplish a comparable effect [5], the processing complexity is intolerable. Besides, two other excellent works, i.e., Cai et al. [4] and Yao et al. [43], have fully considered the temporal-channel-spatial attention and gain exciting outcome. Nonetheless, the above approach does not take into account the nature of the input data as a tensor, which is unavoidable since the training process is accelerated by GPU using parallel computation. In this case, we naturally hope to discover properties that we can use from the relevant theories of tensors to implement the attention mechanism. Drawing inspiration from the tensor CANDECOMP/PARAFAC (CP) decomposition [22], the inverse process of the tensor CP decomposition can be used to calculate the attention map. Therefore, we propose the PFA, which mainly comprises two distinct components, i.e., LPST for the tensor compression and AMC for the attention calculation by tensor fusion. The goal of LPST is to learn parameters that help project the most important information into three 2D tensors. The AMC part uses these three projections to construct an attention map through the inverse process of CP decomposition.
3 Method
In this part, we first introduce the neuron model and the training techniques we adopt. Next, we introduce PFA and its two components, i.e. LPST and AMC, in detail. Finally, a quantitative theoretical analysis on the adoption of connecting factor $r$ is presented.
3.1 Leaky Integrate and Fire Model
In this work, we employ the Leaky Integrate and Fire (LIF) neuron due to its efficiency and simplicity. The LIF neuron accumulates external input in its membrane voltage, which can be described as follows:
$\tau\frac{dV(t)}{dt}=-(V(t)-V_{reset})+I(t),$ (1)
where $\tau$ denotes the time constant, $V(t)$ is the membrane potential at the moment $t$, and $I(t)$ is presynaptic input at $t$ moment. With this formula, the entire model is summarized as:
$\begin{cases}V_{t}^{n}=H_{t-1}^{n}+\frac{1}{\tau}[I_{t-1}^{n}-(H_{t-1}^{n}-V_{reset})]&\\ S_{t}^{n}=\Theta(V_{t}^{n}-V_{thr})\\ H_{t}^{n}=V_{t}^{n}\cdot(1-S_{t}^{n}),\end{cases}$ (2)
where $n$ and $t$ are the layer and moment indices respectively. Besides, $\tau$ represents the time constant, $V$ is membrane potential, $S$ is the spike tensor in binary form, $I$ is the input from the previous layer, $\Theta$ stands for Heaviside step function, $H$ represents the reset behavior after spiking, and $V_{thr}$ means the threshold membrane potential. The LIF model has relatively lower computation complexity and higher representational ability, which now has become a mainstream model in the field of SNN. Here, we use the spatial-temporal backpropagation technique to train our network, aiming to fully make use of the information from spatial and temporal domains. Particularly, the backpropagation on the spatial and temporal domains can be described as follows:
$\begin{cases}\frac{\partial Loss}{\partial S_{i}^{t,l}}=\sum\limits^{n^{l+1}}_{j=1}\frac{\partial Loss}{\partial S_{j}^{t,l+1}}\frac{\partial S_{j}^{t,l+1}}{\partial S_{i}^{t,l+1}}+\frac{\partial Loss}{\partial S_{j}^{t+1,l}}\frac{\partial S_{j}^{t+1,l}}{\partial S_{i}^{t,l}}\\ \frac{\partial Loss}{\partial V_{i}^{t,l}}=\frac{\partial Loss}{\partial S_{i}^{t,l}}\frac{\partial S_{i}^{t,l}}{\partial V_{i}^{t,l}}+\frac{\partial Loss}{\partial S_{i}^{t+1,l}}\frac{\partial S_{i}^{t+1,l}}{\partial V_{i}^{t,l}}.\end{cases}$ (3)
Note that the Heaviside function in (2) is non-differentiable in backpropagation. To address the issue, we employ the following approximation of the Heaviside function in backpropagation for easy calculation:
$ATan(V_{i})=\frac{1}{\pi}arctan[\frac{\pi}{2}\alpha(V_{i}-V_{thr})]+\frac{1}{2},$ (4)
where $V_{i}$ denotes the neuron membrane potential, $V_{thr}$ is the reset membrane potential and $\alpha$ is an adjustment factor to adjust the slope.
3.2 Projected Full-Attention (PFA)
Figure 2: A simple comparison curve of parameter quantity growth among TCJA [48], TA-SNN [42], and PFA. The parameter scale of PFA increases linearly. In this figure, the channel number is fixed at 128.
Figure 3: The detailed workflow of PFA. The input tensor is first sent to Linear Projection of Spike Tensor (LPST) unit to generate three projections and split the three projections into corresponding vectors. In Attention Map Composing (AMC) unit, these vectors are composed into the final attention map through the inverse process of CP decomposition. The attention is fused with the input tensor to obtain the refined tensor by Hadamard product.
In this section, we first introduce the overall idea of PFA. Then we present the details of two crucial components of PFA, i.e., linear projection of spike tensor (LPST) and attention map composing (AMC).
The core of the attention mechanism is to give essential parts of the input data higher weights while ignoring less-related parts to some degree. This is particularly important for SNNs, especially for data collected by DVS cameras. DVS camera data often has more noise, which greatly affects network performance. Thanks to the feature of CP-decomposition, PFA can effectively suppress noise. In our proposed PFA, we first design the LPST module to project the higher-dimensional tensor (generally 3D) to three lower-dimensional tensors, i.e., 2D matrices. By using learnable parameters, LPST can selectively reserve essential information while ignoring irrelevant one. In AMC, we split the matrices into $r$ (called connecting factor here) vectors. Three vectors from three matrices of temporal, channel and spatial dimensions respectively are used to conduct the outer product to compose a part of the final attention map. The final attention map is the sum of these $r$ parts. We fuse this attention map with the origin input to obtain the refined tensor. For an input tensor with the shape of $H\times W\times C\times T$, the overall parameter amount is $C\times r+T\times r+k^{2}\times T\times r$, where $k$ is the convolution kernel size. Note that $r$ is limited to a small fixed value according to our theoretical analysis in Section 3.3, the parameter size grows linearly with $T$ and $C$. Figure 2 shows the parameter growth curve of PFA and compares several other techniques. The workflow of PFA is plotted in Figure 3. In what follows, we will give the details of LPST and AMC.
3.2.1 Linear Projection of Spike Tensor (LPST)
Tensor decomposition is a vital approach to represent a tensor and its latent features. Therefore it is natural for us to utilize tensor decomposition related techniques to explore the data characteristics. A typical technique is the CANDECOMP/PARAFAC (CP) decomposition. Whatever way we adopt, it is essential to generate a “projection” to extract significant components of the input tensor to yield an attention map later.
In this part, we need to construct three linear projections of the input data stream along spatial, channel, and temporal dimensions. In LPST, we apply two fully connected (FC) layers and a convolution layer to project the input tensor for simplicity. Note that here we simply use single-layer FCs and a convolution layer. What we do here is just a simple projection with learnable parameters rather than complex function fitting. In this way, we not only retain the required information, but also reduce the number of parameters.
For temporal dimension, we first squeeze the input tensor ($H\times W\times C\times T$) in the following way:
$Y^{T}_{(c,t)}=\frac{1}{HW}\sum\limits_{i=1}^{H}\sum\limits_{j=1}^{W}X_{(i,j,c,t)},$ (5)
where $Y^{T}_{(c,t)}$ denotes the output matrix, $H$ and $W$ denote the spatial sizes, $X_{(i,j,c,t)}$ denotes the input tensor. After the squeeze operation, the tensor is converted to a matrix in the shape of $C\times T$. Then, we apply an FC layer to reserve $r$ necessary information in each column. A sigmoid activation function is applied after the FC layer. Finally, we obtain a projection matrix in the shape of $r\times T$.
For channel dimension, the operation is similar to that of temporal dimension. We squeeze the input tensor by
$Y^{C}_{(c,t)}=\frac{1}{HW}\sum\limits_{i=1}^{H}\sum\limits_{j=1}^{W}X_{(i,j,c,t)},$ (6)
where $Y^{C}_{(c,t)}$ denotes the output matrix, $C$ denotes the number of channel. Then, we pass the $T\times C$ matrix through an FC layer to acquire the projected matrix in the shape of $r\times T$ with the sigmoid activation function applied.
For spatial dimension, the channel is squeezed by,
$Y^{S}_{(x,y,t)}=\frac{1}{C}\sum\limits_{k=1}^{C}X_{(x,y,k,t)}.$ (7)
Afterward, we employ a convolution to the squeezed tensor, treating $T$ as the input channel and $r$ as the output channel, and then reshape the output tensor from the convolution into a matrix in the shape of $HW\times r$. The sigmoid activation function is applied.
3.2.2 Attention Map Composing (AMC)
Once we obtain the three projections corresponding to three dimensions, a necessary step is to compose an attention map using the three projections. Here, we adopt the reverse process of the tensor CP decomposition. In the CP decomposition, a tensor $\mathcal{X}$ can be represented as:
$\mathcal{X}\approx\sum_{i=1}^{r}v_{si}\circ v_{ci}\circ v_{ti},$ (8)
where $v_{si}$, $v_{ci}$ and $v_{ti}$ denote the $i$-th vectors in the spatial, channel and temporal projection tensors respectively, and $\circ$ is the outer product of vectors. Especially, $r$ is the connecting factor, which can be viewed as the rank concept in the tensor CP decomposition (this process is plotted in Figure 3). It is worth noting that the tensor CP decomposition is just an approximate representation of tensors, which perfectly meets our demands. The accuracy of our representation of the input tensor can be controlled by r. That is the larger $r$ is, the more accurate the representation is, and vice versa, which perfectly realizes the attention mechanism we need. In the experiment part, we discuss the choice of connecting factor $r$ and its effect in details.
3.3 Quantitative Theoretical Analysis
Figure 4: A toy example of the effect of different ranks on the approximation outcome. error is measured under $\ell_{2}$ norm. The norm starts to cover when the rank exceeds 30.
In this section, we first introduce an approach to estimate the rank of the input spike tensor. Following such analysis, we propose a principle to select the proper $r$ value, which is supported by our experiments.
To better understand the feature of PFA and facilitate later application, we follow the work by Acar et al. [1] to investigate the rank of the input data. Given that computing the CP decomposition of a tensor is recognized as an NP-hard problem [13; 22], direct calculation from the tensor becomes infeasible. Therefore, to address this challenge, we adopt a gradient descent technique to approximate the original tensor. We manually set the rank to multiple values, and consider the value with the best approximation as the rank of the tensor. To be more specific, an element of the tensor can be presented as:
$x_{ijk}\approx\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r}.$ (9)
We treat it as an optimization problem, in which $a_{i,r},b_{j,r},c_{k,r}$ are considered as parameters. Then we optimize the $Loss$ in the following expression using the gradient descent approach:
$Loss=\frac{1}{2}\sum_{(i,j,k)\in\Omega}(x_{i,j,k}-\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r})^{2}.$ (10)
Find the first order partial derivative of $Loss$ about the parameter ($a_{i,r}$ as an example here), namely:
$\frac{\partial Loss}{\partial a_{i,r}}=-\frac{1}{2}\cdot 2\cdot(x_{i,j,k}-\sum\limits_{r=1}^{R}a_{i,r}b_{j,r}c_{k,r})\cdot b_{j,r}c_{k,r}.$ (11)
Lastly, apply the gradient descent method to execute optimization in the following way:
$a_{i,r}^{\prime}\leftarrow a_{i,r}-\mu\cdot\frac{\partial Loss}{\partial a_{i,r}},$ (12)
where $\mu$ denotes the step size of the gradient descent. We take a piece of data from the dataset CIFAR10DVS as a toy example and discover its rank. In this case, we adopt $\mu=0.0001$ and iterate the process 1000 times to get the final optimized parameters. We utilize $\ell_{2}$ norm as the metric to judge the approximation effect. As shown in Figure 4, the error reduces as the rank increases, and starts to converge when the rank exceeds 30. This toy example shows that the rank of the input tensor is around 30, indicating that we need a rank exceeding 30 to precisely represent it using the CP decomposition technique. It should be noted that for the attention mechanism, representing the original tensor too accurately is not conducive to improving network performance. The core idea of the attention mechanism is to focus on the important parts while ignoring the secondaries. Representing raw tensors too precisely is counterproductive. Therefore, a good choice for $r$ should not be too large. This is a very useful conclusion. On the one hand, it tells us the range of finding $r$. On the other hand, it also ensures that PFA can obtain good results without excessive calculation.
As shown in our experiment (Figure 6), though the rank of the attention map differs on dynamic datasets like CIFAR10DVS and static datasets like CIFAR100, the best option for $r$ is not very big. Further observing the experimental results, we find that when the value of $r$ exceeds a certain threshold (approximately $\frac{T}{2}$, $T$ is the time step), the larger $r$ is, the worse the effect will be. The above analysis of the experimental results proves the correctness of our conclusion. Considering the above analysis, we propose a principle for selecting $r$: For dynamic datasets, the search for the optimal value of $r$ is performed within the range of no more than $T$. Specifically, we focus on identifying the best $r$ value around $\frac{T}{2}$, considering the varying content across different time steps. In the case of static datasets, the search for the optimal $r$ value begins from 1 and extends until the best value is determined. Since static datasets involve duplicated frames, this search aims to find the most suitable $r$ value for the low-rank condition.
4 Experiment
We evaluate the classification performance of our PFA on two static datasets: CIFAR10 [24], CIFAR100 [24] and two dynamic datasets: CIFAR10DVS [26], NCaltech-101 [32]. Details of the datasets, network architecture, data augmentation, loss function, and pre-process procedure is introduced in this section. Extensive experiment results are also presented.
Dataset Learning Rate Epoch Batch Size
CIFAR10 0.1 200 128
CIFAR100 0.1 200 64
CIFAR10DVS 0.0001 200 32
NCaltech-101 0.0001 200 32
Table 1: Hyper-parameter settings of PFA.
4.1 Datasets and Training Details
4.1.1 Datasets
We have conducted experiments on both static datasets and dynamic datasets for object classification. We train and test PFA on a workstation equipped with one RTX 3090. For different datasets, the hyper-parameters are listed in Table 1. The summaries of datasets and augmentation involved in the experiment are listed below.
CIFAR 10/100 consist of 50k training images and 10k testing images with the size of 32 × 32 [24]. We use ResNet-19 for both CIFAR10 and CIFAR100. Random horizontal flips and crops are applied to the training images for augmentation. Moreover, Cutout is also used for augmentation which is the same as the Ref [8]
CIFAR10-DVS converts 10,000 frame-based images of 1010 classes into event streams with the dynamic vision sensor. Since the CIFAR10DVS dataset [26] does not divide training and testing sets, we split the dataset into 9k training images and 1k test images and reduced the spatial resolution from $128\times 128$ to $48\times 48$ [8]. We use VGGSNN for neuromorphic datasets CIFAR10DVS. In each frame, horizontal flipping and mixup are adopted, where the probability of Flipping is set to 0.5. Then, we randomly select one augmentation among Rolling, Rotation, Cutout, and Shear, where the random Rolling range is 5 pixels, and the degree of Rotation is sampled from the uniform distribution, which is the same as the Ref [8].
NCALTECH-101 is also converted from the original version of Caltech-101 [32] with a slight change in object classes to avoid confusion. The NCaltech-101 consists of 100 object classes plus one background class. We apply the 9: 1 train-test split as CIFAR10DVS. And the VGGSNN is applied to neuromorphic datasets NCALTECH-101. Moreover, the augmentation is identical to CIFAR10DVS.
4.1.2 Loss function
We used the cross-entropy loss for the neuromorphic datasets CIFAR10DVS and NCALTECH-101. And in order to get a better performance in static datasets (CIFAR 10 / 100), we used a Temporal Efficient Training [8] loss function. It can be concluded as follows:
$\mathcal{L}_{\mathrm{TET}}=\frac{1-\lambda}{T}\cdot\sum_{t=1}^{T}\mathcal{L}_{\mathrm{CE}}[\boldsymbol{O}(t),\boldsymbol{y}]+\frac{\lambda}{T}\cdot\sum_{t=1}^{T}\mathcal{L}_{\mathrm{MSE}}(\mathbf{O}(t),\phi),$ (13)
where $T$ is the total time steps, $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{MSE}}$ denote the cross-entropy loss and mean square error. And $\boldsymbol{y}$ represents the target label. Moreover, $\mathbf{O}(t)$ and $\phi$ are the output and a constant used to regularize the membrane potential distribution. Additionally, we set $\phi=V_{th}$ in our experiments. In practice, we use a hyperparameter $\lambda$ to adjust the proportion of the regular term.
Figure 5: Schematic diagram of the network structure. On the left is the convolutional layer of the VGG network, and on the right is the convolutional layer of Resnet.
4.1.3 Network Architectures
To evaluate PFA, we integrate it into VGG-type [37] and ResNet [14] architectures, which are previously utilized in [12; 48; 8; 43]. Figure 5 presents the basic blocks involved in VGG and ResNet architectures. Five PFAs are inserted into a VGG-type network after each average pooling operation. In ResNet architecture, PFA is inserted before the residual connection in the last two layers. When splitting the event flow data into frames, we simply integrate the input in the time domain. Particularly, our work is based on the software framework SpikingJelly [10]. Taking notice that frames in CIFAR10-DVS and NCaltech-101 are entirely occupied by their content, we utilized a pooling layer as the first layer of the network to reduce memory consumption.
4.2 Comparison with Existing SOTA Works
To verify the effectiveness of our PFA, we compare it with several current best SNNs, including hybrid training and spike-based backpropagation. The results for the static image data and the dynamic data classification task are listed in Table 2. The results for the static image dataset are at time steps 2, 4, and 6.
For CIFAR10 and CIFAR100, PFA achieves higher accuracy than previous works. In particular, our PFA achieves superior performance on CIFAR10 and CIFAR100, with 1.2% and 4.4% improvements over the TET [8] method, respectively, using the same network structure and time steps. Moreover, our PFA surpasses the DSR [30] method while requiring 10$\times$ fewer time steps. Additionally, we demonstrate that our attention-based SNN outperforms the Spiking-ViT [47] backbone.
On the dynamic dataset CIFAR10DVS, our method outperforms the previous works with binary spikes by 3.1%, while using fewer time-steps. For the NCaltech101 dataset, we achieved 80.5% top-1 accuracy, achieving a performance far beyond previous work [48]. To further illustrate the advantages of our PFA, we show the comparison with previous works at different time steps. As shown in Figure 1, we compared PFA with Spike-based BP SNNs. Specifically, our PFA shows a significant advantage at shorter simulation lengths due to its more vital representation ability.
Dataset Method Type Architecture Time steps Accuracy CIFAR10/CIFAR100 SIDP [35]ICLR-2019 Hybrid training ResNet-20/VGGSNN 250/125 95.4/76.4 Diet [34]TNNLS-2021 Spike-based BP ResNet-20 10/5 92.5/64.0 tdBN [46]AAAI-2021 Spike-based BP ResNet-19 6 93.1/71.1 Dspike [27]NeurIPS-2021 Spike-based BP ResNet-18 6 94.2/74.2 NAS-SNN [21]ECCV-2022 Spike-based BP NAS 5 92.7/73.0 TET [8]ICLR-2022 Spike-based BP ResNet-19 6 94.5/74.7 DSR [30]CVPR-2022 Spike-based BP ResNet-19 20 95.4/78.5 Spikformer [47]ICLR-2023 Spike-based BP Spiking-ViT 4 95.5/78.2 PFA Spike-based BP ResNet-19 6 95.7/79.1 4 95.7/78.1 2 95.6/76.7 CIFAR10DVS LIAF-Net [40]TNNLS-2021 Spike-based BP LIAF-Net 10 70.4 tdBN [46]AAAI-2021 Spike-based BP ResNet-19 10 67.8 TA-SNN [42]ICCV-2021 Spike-based BP LIAF-Net 10 72.0 Dspike [27]NeurIPS-2021 Spike-based BP VGGSNN 10 75.4 TET [8]ICLR-2022 Spike-based BP VGGSNN 10 83.2 DSR [30]CVPR-2022 Spike-based BP VGGSNN 10 77.3 TCJA [48]arXiv-2022 Spike-based BP VGGSNN 10 80.7 Spikformer [47]ICLR-2023 Spike-based BP Spiking-ViT 16 80.9 PFA Spike-based BP VGGSNN 14 84.0 NCaltech-101 SALT [20]Neural Netw-2021 Spike-based BP VGGSNN 20 55.0 TCJA [48]arXiv-2022 Spike-based BP VGGSNN 14 78.5 PFA Spike-based BP VGGSNN 14 80.5
Table 2: Compare with existing works. Our method improves network performance across all tasks.
4.3 Discussion on $r$
Refer to caption CIFAR10
Refer to caption CIFAR100
Refer to caption CIFAR10DVS
Refer to caption NCaltech101
Figure 6: Accuracy on four tested datasets with different $r$. PFA achieves best performance when $r=3,1,8,8$ correspondingly.
Figure 6 demonstrates the performance with different $r$. The outcome reveals that for dynamic datasets, the best rank option falls around $\frac{T}{2}$ while for static datasets, the best value of $r$ (also known as connecting factor) is very small. Dynamic datasets exhibit significant variations in content across different time steps. Conversely, static datasets involve the duplication of frames, where each frame is replicated $T$ times as input to the network. Consequently, this duplication results in the attention map being subjected to a low-rank condition. Considering a comprehensive assessment of both experimental outcomes and theoretical insights, our conclusion asserts that the precision of the attention map in faithfully mirroring the input tensor is unnecessary. Such fidelity does not inherently facilitate the effective concentration on pivotal components. Furthermore, our findings unveil an intriguing phenomenon: even in scenarios where SNNs exhibit considerable sparsity, the rank of the attention map can still remain markedly low. This revelation implies the presence of untapped optimization potential within the realm of SNN sparsity.
4.4 Ablation Study
In this section, we conduct ablation experiments from both global and local perspectives to verify the reasonableness of PFA. We first try adding different numbers of PFA layers to the network to see how PFA improves classification accuracy. Then we set the three projections obtained inside the PFA to a constant matrices with each element assigned to 1 respectively to eliminate the attention on this dimension to explore the impact of different dimensions of attention on the classification results.
We insert different numbers of PFAs into the VGG network and obtain the results in Figure 7 (left). In this experiment, $r$ is fixed at 8. We can observe that as more and more PFA layers are inserted, the accuracy of the network is getting higher and higher. Even inserting only one layer can improve the accuracy of the network.
In order to explore how attention in temporal-channel-spatial dimensions affects classification accuracy, we conduct an ablation by fixing corresponding projection matrix obtained by LPST to a all-one tensor. As shown in Figure 7 (right), whichever dimension is ablated will adversely affect the results. Ablating temproal and ablating channel have similar effects. This means that PFA treats temporal and channel dimensions equally. However, when we ablate the spatial dimension, the accuracy of the network drops significantly (79.60%) but still outperforms the baseline (77.90%). Whichever dimension we ablate, the performance drops. This observation highlights that PFA effectively directs attention to distinct regions across multiple dimensions, aligning with the intended purpose of attention modules. Specifically, PFA exhibits attention distribution across all three dimensions, albeit with varying weights. Notably, the spatial dimension emerges as the most significant, followed by the time and channel dimensions, in terms of attention importance.
Figure 7: Results of ablation experiments. The left one is ablation study on PFA layers. The x-axis represents the amount of PFA inserted. The right one is the result of ablation study on temporal-channel-spatial ablation. Full denotes the original PFA without any modification as a reference. w.o. T/C/S denotes setting the temporal/channel/spatial projection to a constant matrix with each element assigned to 1.
4.5 Attention Visualization
In order to visually represent the attention distribution of the PFA module, we have generated an attention map in PFA and depicted it in Figure 8. The attention maps corresponding to the temporal and channel dimensions in the figure are obtained from the deep layer of the network. This choice was made because deep layers tend to provide a more accurate reflection of the attention’s role. Conversely, the spatial attention map is obtained from the input layer of the network, where the shallow graphic features are more distinct and conspicuous, facilitating analysis and observation. Figure 8 also illustrates the distinctive attention distributions in the temporal, channel, and spatial dimensions achieved by PFA. This finding aligns with the conclusions drawn from our ablation study, further validating our earlier assertions.
5 Conclusions
In this paper, we introduce the novel approach named PFA, which combines learning and generation of a temporal-channel-spatial attention map through CP decomposition. Specifically, we propose the LPST to learn a projection of the input tensor that can extract essential information for AMC to compose the attention map. In the AMC, we utilize the reverse process of CP decomposition to generate the attention map. After analyzing the rank of the raw input tensor and the best option for connecting factor $r$, we propose a principle to search for the best $r$. Remarkably, our findings suggest that even a low rank can lead to outstanding performance. Regardless the form of the dataset, dynamic or static, PFA achieves the SOTA performance. Additional ablation studies prove the reasonableness of PFA while visualization of PFA attention distribution helps us intuitively understand the significance effect of PFA.
Figure 8: Attention visualization on CIFAR10DVS dataset. Temporal and channel parts are taken from the deep layer of VGGSNN. The spatial layer visualizations are extracted from the shallow layer of the network due to the inherent complexity of the information contained in the deeper layers, which may impede human comprehension and hinder the interpretability of our analysis.